OpenStudy (anonymous):

derite of 2 / sqareroot of T^3 is this 2t^-1/3

OpenStudy (anonymous):

2/(squareroot (t^3)) = 2/(t^(3/2) = 2t^(-3/2) Taking the square root is the same as raising to the 1/2 power.

OpenStudy (anonymous):

OK I THOUGT FOR SOME REASON IT WAS 2T^-1/3

OpenStudy (anonymous):

are you asking for the derivative?

OpenStudy (anonymous):

because then its -3/2(2t)^-5/2

OpenStudy (anonymous):

which is is -3t^-5/2

OpenStudy (anonymous):

YES DERIVATE OF 2 DIVIDED SQTROOT OF T^3

OpenStudy (anonymous):

(d/dt) [2t^(-3/2)] = -3t^(-5/2) Just rewrite with a negative exponent and use the power rule. Alternatively, you could use the quotient rule, but this is the easiest way when you don't have any exponents in the numerator.

OpenStudy (anonymous):

*variables, not exponents